Fermi National Laboratory
 

Operation Summary for January 20, 2010 - January 22, 2010

 

Wednesday Jan. 20

The day shift began with Operations monitoring store 7520 (112.4E30) and the stash (195.3E10), with the Antiproton Source (Pbar) stacking (10.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and MTest T977 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, at 1 PM, Pbar Personnel accessed their rings to investigate a LCW leak. They found a leak on an Accumulator quadrupole magnet. Linac personnel flushed the H- Source water resistor.

During the evening shift, Operations disabled the MTest beam at 4:30 PM when a P2 transfer line correction magnet power supply (VTF11) failed. EE Support personnel had the supply repaired by 7:15 PM. An AP1 transfer line quadrupole power supply failed to switch to its reverse proton setting at 7:45 PM. Operators replaced a control card and transfers to Recycler continued.

Thursday Jan. 21

The midnight shift began with Operations monitoring store 7520 (21.03E30) and the stash (407.1E10), with Pbar stacking (19.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7520 at 12:19 AM and then put the TeV into a HEP shot setup. Operations established store 7521 at 2 AM with an average initial luminosity (AIL) of 285.03E30. Operations established beam to MTest at 6:35 AM after operators moved the MI septa back to its correct position.

During the day shift, Operations terminated store 7521 at 2:20 PM and TeV proton-only studies began.

During the evening shift, TeV studies concluded at 6:35 PM and Operations put the TeV into shot setup at 7:05 PM. Operations established store 7523 at 8:30 PM with an AIL of 255.5E30.

Friday Jan. 22

The midnight shift began with Operations monitoring store 7523 (139.98E30) and the stash (145.1E10), with Pbar stacking (27.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, store 7523 aborted at 3:30 AM due to a failed shunt interlock chassis at TeV sector C0; there was no quench. EE Support personnel came in and replaced the chassis and the auxiliary power supply unit. Operations established beam to MTest at 4:20 AM. Operations began the TeV turn on at 5:45 AM. At the end of the shift, a failed power supply in the CDF detector required an access into the collision hall to make repairs.

The Plans for Friday and the Weekend

The plans for today are for TeV personnel to investigate an emittance problem and then for the rest of the weekend to resume normal stacking, stashing, and storing.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/03/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory